Heat Physics

Occasionally, huge icebergs are found floating on the ocean's currents. Suppose one such iceberg is 147 km long, 30.4 km wide, and 204 m thick. (a) How much heat injoules would be required to melt this iceberg (assumed to be at 0 °C) into liquid water at 0 °C? The density of ice is 917 kg/m3. (b) The annual energy consumption bythe United States in 1994 was 9.3 x 1019 J. If this energy were delivered to the iceberg every year, how many years would it take before the ice melted?
Answer #1
let volume of ice,v=147000*30400*204 m^3
now,weight,m=917*v

1.Q=mL L=latent heat of fusion of water=334kj/kg
=>Q=147000*30400*204*917*334=2792138057856*10^8 J=27.921*10^19 J

2.years it would supply energy=Q/9.3*10^19
=27.921*10^19/9.3*10^19
=3 years

Hope my answer helps you.thank you.
answered by: lanita